Huang Automotive is presently operating at 75% of capacity. The company recently received an offer from a Korean truck manufacturer to purchase 26,000 units of a power steering system component for $200 per unit. Peter Wu, vice-president of sales, notes that although there will be an additional $2.25 shipping cost for each component, he thinks that accepting the order will get the company's "foot in the door" of an expanding international market.

To determine variable and fixed costs, Huang's accountant used the high-low method with the following production and cost information for the last two years:

192,000 units     230,000 units    
Direct material costs $16,704,000      $20,010,000     
Direct labor costs 5,760,000      6,900,000     
Overhead costs 24,752,000      26,880,000     
Selling and administrative costs 12,304,000      12,760,000     
Total costs $59,520,000      $66,550,000     
Total costs per unit $310.00      $289.35     



T.J. Chan, vice-president of engineering, feels that any new market should first show its profitability and that the $200 per unit offer is not only below the regular $270 selling price, but it's below the unit cost of the component. She also points out that there will be additional setup costs of $260,000 and that Huang will have to lease some special equipment for $200,000.

REQUIRED [6 tries]
1. Using the high-low method to determine cost behavior, what would the expected profit be on the special order (use a negative sign for a loss)?
Answer #1
Units 192,000 230,000
Direct material 16,704,000 20,010,000
Direct Labor 5,760,000 6,900,000
Overhead costs 24,752,000 26,880,000
Selling and Admin cost 12,304,000 12,760,000
Total cost 59,520,000 66,550,000
Using high low method, variable cost per unit = Difference in Cost/Difference in Volume
Overhead 56 per unit
Selling and Admin Cost 12 per unit
Direct material and direct labor are variable costs
Since there is spare capacity, only incremental costs are relevant for the special order
Calculation of profit from special order:
Revenue 5,200,000
Less: costs
Direct material 2,262,000
Direct Labor 780,000
Overhead costs 1,456,000
Selling and Admin cost 312,000
Shipping cost 58,500
Setup costs 260,000
Lease cost 200,000
Total cost 5,328,500
Profit from special order -128,500

в 1 2 Units 192000 230000 3 Direct material 16704000 =B3*C2/B2 4 Direct Labor 5760000 =B4*C2/B2 5 Overhead costs 24752000 268